1.the amount of charge possessed by 1kg of electrons?
2.A partice os mass m and carrying charge -q1 is moving around a charge +qalong a circular path of radius r.find period of revolution of the charge -q1.
3.The electric field at 2R from the centre of a uniformly charged non-conducting sphere of radius R is E.The electric field at a distance of R/2 from the centre will be?(the answer was given as 2E..but i’ve studied that the electric field at a point inside a uniformly charged sphere is 0?)
 

Grade:

2 Answers

Abhishek Kumar
9 years ago
\therefore1
1e has charge of 1.602 X 10-19 C
So, 9.109 X 10-31 kg has 1.602 X 10-19 C
So, 1 Kg has = 1.602 X 10-19/(9.109 X 10-31) C

2
Kq1q2/r2 = mv2/r Here we need one more equation as there are two variables and only one equation. Or, assumne mvr= nh/2\pi and then solve.
T= 2\pir/v

3
Electric field inside a uniformly charged dielectric sphere is NOT zero. It is zero inside a metallic sphere or any metal.
Electric field inside a uniformly charged dielectric sphere is = kqr/R3
Kushagra Madhukar
3 years ago
Dear student,
Please find the attached answer to your problem.
 
1
1e has charge of 1.602 X 10-19 C
So, 9.109 X 10-31 kg has 1.602 X 10-19 C
So, 1 Kg has = 1.602 X 10-19/(9.109 X 10-31) C
2
Kq1q2/r2 = mv2/r Here we need one more equation as there are two variables and only one equation. Or, assumne mvr=
nh/2π and then solve.
T = 2πr/v
3
Electric field inside a uniformly charged dielectric sphere is NOT zero. It is zero inside a metallic sphere or any metal. Electric field inside a uniformly charged dielectric sphere is = kqr/R3
 
Hope it helps.
Thanks and regards,
Kushagra